OpenStudy (anonymous):

Given that f(x)=-2h(x)+(2x/h(x)); h(4)=4, h'(4)=-2. Find f'(4)

OpenStudy (anonymous):

well their is: a)13/2 b)7/2 c)-5/2 d)11/2 e)17/2

OpenStudy (anonymous):

I tried solving for the derivative of f(x) but I must be doing something wrong or the awnser is not there

myininaya (myininaya):

So what did you get for f'?

myininaya (myininaya):

I will check it.

OpenStudy (anonymous):

\[f'(x) = h(x) + -2h'(x)+ (\frac{ 2h(x)-(2x)h(x) }{ h(x)^2 }\]

myininaya (myininaya):

Also the answer is there. h(x) shouldn't be there. you are missing h' in your quotient

myininaya (myininaya):

(-2h)'=-2(h)'=-2h'

OpenStudy (anonymous):

oh ok

myininaya (myininaya):

(2x/h)'=[(2x)'h-2x(h)']/h^2

OpenStudy (anonymous):

so the awnser should be 19/2

myininaya (myininaya):

hmm... that isn't what I got...let me check my result.

myininaya (myininaya):

Oh you are still assuming h is there aren't you?

myininaya (myininaya):

(-2h)'=-2(h)'=-2h' not h-2h'

OpenStudy (anonymous):

So how does it apply

myininaya (myininaya):

what do you mean?

OpenStudy (anonymous):

well I am plugging in everything where \[h(4)+(-2h'(4)) + \frac{ (2)(h(4))-(2x)(h'(4)) }{ h(4)^2 } = 4 +(-2(-2)) + \frac{ (2)(4)-(2(4))(-2) }{4^2 }= f'(4)\]

OpenStudy (anonymous):

correct?

myininaya (myininaya):

Well see I keep telling you (-2h)'=-2h' not h-2h'

myininaya (myininaya):

That first term you are putting in shouldn't be there.

OpenStudy (anonymous):

Could you show me what you plugged in so I understand

myininaya (myininaya):

So you don't understand why (-2h)' is -2h' and not h-2h'?

OpenStudy (anonymous):

No

myininaya (myininaya):

I'm liking I'm using the constant multiple rule.

myininaya (myininaya):

(cf)'=cf' Let k(x)=cf(x) To find k'(x) we use the formal definition of derivative: \[k'(x)=\lim_{\Delta x \rightarrow 0} \frac{k(x+\Delta x)-k(x)}{\Delta x}=\lim_{\Delta x \rightarrow 0} \frac{cf(x+\Delta x)-cf(x)}{\Delta x}\] Note: By the distributive property ab+ac=a(b+c) \[=\lim_{\Delta x \rightarrow 0} c \frac{f(x+\Delta x)-f(x)}{\Delta x} \] Note: By limit properties, we can do lim c g(x) = c lim g(x) \[=c \lim_{\Delta x \rightarrow 0} \frac{f(x+\Delta x)-f(x)}{\Delta x} \] Note: By definition of derivative that above is cf'(x).

myininaya (myininaya):

The constant multiple rule: k=cf k'=cf'

myininaya (myininaya):

So say k=ch then k'=ch' That c means constant The constant in front of h in your problem is -2 so k=-2h then k'=-2h'

OpenStudy (anonymous):

So when I plug in for -2h' when h(4)=4 and h'(4)=-2

myininaya (myininaya):

Yep

myininaya (myininaya):

-2h'(x) x is 4. -2h'(4) -2h'(4)=-2(-2)=4 Now simplify your quotient and add.

OpenStudy (anonymous):

Does this mean -2h'= -2(-2) or -2(4)

OpenStudy (anonymous):

ok

myininaya (myininaya):

h' is h'(x)

myininaya (myininaya):

[-2h(x)] ' =-2 [ h(x)] ' =-2 h'(x) This has been what I have been saying but you wrote it as h(x)-2h'(x)

myininaya (myininaya):

\[\cancel{h(4)}+(-2h'(4)) + \frac{ (2)(h(4))-(2x)(h'(4)) }{ h(4)^2 } = \cancel{4} +(-2(-2)) + \frac{ (2)(4)-(2(4))(-2) }{4^2 }= f'(4)\]

myininaya (myininaya):

This is what you wrote earlier I'm telling you (-2h)'=-2h' not h-2h' do you understand now?

OpenStudy (anonymous):

Oh ok Now I get what you mean by this

myininaya (myininaya):

I put the marks through the bad parts.

myininaya (myininaya):

so you understand the constant multiple rule?

OpenStudy (anonymous):

Yes I do now understand the multiple rule and the awnser is 11/2

myininaya (myininaya):

That is right.

OpenStudy (anonymous):

I just didnt understand it conceptually but now that I see it plugged in and used I do now

myininaya (myininaya):

Ok. Neatness.

OpenStudy (anonymous):

Thank you so much

myininaya (myininaya):

Np.